The television show Found has been successful for many years. That show recently had a share of 19, meaning that among the TV sets in use, 19% were tuned to Found. Assume that an advertiser wants to verify that 19% share value by conducting its own survey, and a pilot survey begins with 14 households have TV sets in use at the time of a Found broadcast.
Find the probability that none of the households are tuned to Found.
P(none) =
(round answer to 4 decimal places)
Find the probability that at least one household is tuned to Found.
P(at least one) =
(round answer to 4 decimal places)
Find the probability that at most one household is tuned to Found.
P(at most one) =
(round answer to 4 decimal places)
If at most one household is tuned to Found, does it appear that the 19% share value is wrong? (Hint: Is the occurrence of at most one household tuned to Found unusual?)
yes, it is wrong
no, it is not wrong

Answer by stanbon(75887)

The television show Found has been successful for many years. That show recently had a share of 19, meaning that among the TV sets in use, 19% were tuned to Found. Assume that an advertiser wants to verify that 19% share value by conducting its own survey, and a pilot survey begins with 14 households have TV sets in use at the time of a Found broadcast.
P(watch) = 0.19 ; p(don't watch) = 0.81
-------------------------
Find the probability that none of the households are tuned to Found.
P(none) = (0.81)^14 = 0.0523
(round answer to 4 decimal places)
------------------------------------------
Find the probability that at least one household is tuned to Found.
P(at least one) = 1-P(none) = 0.9477
(round answer to 4 decimal places)
------------------------------------------
Find the probability that at most one household is tuned to Found.
P(at most one) = binomcdf(14,0.19,1) = 0.2242
(round answer to 4 decimal places)
-----------
If at most one household is tuned to Found, does it appear that the 19% share value is wrong? (Hint: Is the occurrence of at most one household tuned to Found unusual?)
Not unusual.
----
Cheers,
Stan H.
-----------
yes, it is wrong
no, it is not wrong